# TPK: Trustworthy Trajectory Prediction Integrating Prior Knowledge For Interpretability and Kinematic Feasibility

### v3

@@ Abstract
Index Terms:
Trajectory Prediction, Motion Planning
@@ III-B2 Kinematic Prior Integration
We derive the control input limits based on physical constraints and empirical observations of vehicle and pedestrian behavior. For vehicles, we adopt the well-established limits of [8,8][-8,8][ - 8 , 8 ] m/s2 for acceleration and [0.3,0.3][-0.3,0.3][ - 0.3 , 0.3 ] 1/m for curvature [5, 8]. For pedestrians, as prior studies lack specific limits, we draw on research into human walking and acceleration behavior [30, 31, 32] to define appropriate constraints. These studies suggest that a reasonable maximum acceleration for pedestrians is [8,8][-8,8][ - 8 , 8 ] m/s2 while their speed is limited to [0,10][0,10][ 0 , 10 ] m/s.
@@ V-B1 Infeasible Motions
Model
Kinematic
Layer
Class-Specific
Interaction
Layers
Interaction
Prior
Integration
All Agents Vehicle Pedestrian Cyclist
Brier-
minFDE6
mAP ($\uparrow$) minADE6 minFDE6 minADE6 minFDE6 minADE6 minFDE6 minADE6 minFDE6
@@
Model
Brier-
minFDE6
mAP ($\uparrow$) minFDE6 minFDE1 MR2 minADE6 minADE1
Infeasible
Steps
Infeasible
Predictions
$\rho\,$(minADE6, $\Delta\alpha$) ($\uparrow$)
@@ VI Conclusion
Copyright Notice
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.

### v4

@@ III-B2 Kinematic Prior Integration
We derive the control input limits based on physical constraints and empirical observations of vehicle and pedestrian behavior. For vehicles, we adopt the well-established limits of [8,8][-8,8] m/s2 for acceleration and [0.3,0.3][-0.3,0.3] 1/m for curvature [6, 13]. For pedestrians, as prior studies lack specific limits, we draw on research into human walking and acceleration behavior [32, 15, 11] to define appropriate constraints. These studies suggest that a reasonable maximum acceleration for pedestrians is [8,8][-8,8] m/s2 while their speed is limited to $[0,10]$ m/s.
@@ V-B1 Infeasible Motions
Model Kinematic Layer Class-Specific Interaction Layers Interaction Prior Integration All Agents Vehicle Pedestrian Cyclist
Brier- minFDE6 mAP ($\uparrow$) minADE6 minFDE6 minADE6 minFDE6 minADE6 minFDE6 minADE6 minFDE6
@@
Model Brier- minFDE6 mAP ($\uparrow$) minFDE6 minFDE1 MR2 minADE6 minADE1 Infeasible Steps Infeasible Predictions $\rho\,$(minADE6, $\Delta\alpha$) ($\uparrow$)